OpenStudy (anonymous):

What is the slope of -2x+3y=6

Nnesha (nnesha):

first solve for y so add 2x both side

OpenStudy (anonymous):

1. subject y : 3y= 6 +2X y= 2 + 2x/3 2. the co-efficient with x is slope : 2/3